if the perimeter is 100 meters and a the length is 14m more than the width what is the length and the width?
Answer by cleomenius(959)

x = w
x + 14 = L
2x + 2x + 28 = 100
4x = 72
x = 18 meters the width.
32 = meters the length
2(18) + 2(32) = 100 meters
Cleomenius.